You are considering how to invest part of your retirement savings.You have decided to put

$100,000 into three?stocks:

54%

of the money in GoldFinger? (currently

$16?/share),

11 %

of the money in Moosehead? (currently

$88?/share),

and the remainder in Venture Associates? (currently

$7?/share).

Suppose GoldFinger stock goes up to

$39?/share,

Moosehead stock drops to

$63?/share,

and Venture Associates stock

risesrises

to

$12

per share.
Answer #1

Money invested in Goldfinger = 54% of $100,000 = $54,000

No. of stocks of Goldfinger purchased = $54,000/$16 = 3375 shares.

Money invested in Moosehead = 11% of $100,000 = $11,000.

No. of stocks of Moosehead = $11,000/$88 = 125 shares

Money invested in Venture Associates = 100,000-(54,000-11,000) = $35,000

No. of stocks of Venture Associates = $35,000/$7 = 5000 shares

New value of portfolio = (3375 shares*$39) + (125 shares*$63) +(5000 shares*$12)

= $199,500.

1. Thus portfolio value after all changes in stock prices are accounted for = $199,500

2. % change in portfolio = (199500-100000)/100000 = 99.50%

3. Weight of each stock in the portfolio:

Weight of Goldfinger = (3375 shares*$39)/199500 = 65.98%

Weight of Moosehead = (125 shares*$63)/199500 = 3.95%

Weight of Venture Associates = (5000 shares*$12)/199500 = 30.08%